Two numbers p and q have a sum of 27 Two numbers p and q have a sum of 27. Find the numbers
Answer by Cromlix(4381)

Hi there,
Two numbers p and q have a sum of 27
p + q = 27......(1)
and a difference of 9.
p - q = 9.......(2)
...........
p + q = 27......(1)
p - q = 9.......(2)
Add Eqs(1) + (2)
2p = 36
p = 18.
Substitute p = 18 into
Eq (1)
p + q = 27......(1)
18 + q = 27
q = 27 - 18
q = 9.
.......
p = 18
q = 9
Hope this helps :-)